OpenStudy (anonymous):

Using complete the square, if x^2-6x-3=(x-h)^2+k. What is the value of k

hero (hero):

x^2-6x = 3 x^2 - 6x + 9 = 3 + 9 (x-3)^2 = 12 (x-3)^2-12 = (x-h)^2+k k = -12